Question 03

Write (and draw the graph of) a function which is

(a) Continuous on all points except at x = 1.

(b) Differentiable on all points except at x = 1.

(c) Non-differentiable at five points x = 1, x = 2, x = 3, x = 4 and x = 5.


1
Expert's answer
2022-01-18T10:36:32-0500

a,b)

"y=\frac{1}{x-1}"




c)

"y=\frac{1}{x-1}+\frac{1}{x-2}+\frac{1}{x-3}+\frac{1}{x-4}+\frac{1}{x-5}"